Two forces act on a point object as follows 100N at 170°and 100N at 50°.find their resultant
1
Expert's answer
2019-10-08T10:03:31-0400

Angle between the two forces are:

"170\degree-50\degree=120\degree"

So ,

Resultant force ,

"F=\sqrt{F_1^2+F_2^2+2F_1F_2\cos\theta}"

Now,

"F_1=F_2=100\ N"

And

"\theta=120\degree"

So,

"F=\sqrt{100^2+100^2+2×100×100×\cos120\degree}\\=100\ N"

Direction:

Since both forces are equal so the direction of resultant force will be exactly midway between these two forces

So angle will be "\frac{(170+50)}{2}=110\degree"